# Using matrix-product states for time-series machine learning

### v1

@@ III.2.1 Synthetic time series
Here, we investigate the ability of MPSTime to impute missing values from time-series data distributions of incrementally increasing complexity. Simulated instances were randomly split between training (60%) and testing (40%) sets. We varied the complexity of the dataset by changing the trend and period of the sinusoid, allowing us to systematically assess our method’s imputation performance across a range of challenging settings. Full details of the time-series model we used and its corresponding parameter values are in Sec. III.1.1.
For our simplest imputation setting, we examined two levels of additive Gaussian noise: (i) low noise ($\sigma=0.1$); and (ii) moderate noise ($\sigma=0.2$), where $\sigma$ denotes the standard deviation of the Gaussian noise distribution. In both cases, we held the period $\tau$ and trend $m$ fixed while allowing the phase to vary, which, for each time series, was drawn randomly from a uniform distribution, $\psi\sim U(0,2\pi)$. We then trained a $d=12$ MPS with three bond dimensions ($\chi_{\rm max}=20,30,40$) for ten sweeps on each of the training sets. For every test instance, we removed a central block of consecutive time-series points (with size determined by the percentage of total samples missing), maintaining approximately equal numbers of measured points on either side of the missing region.
Our first key result is presented in Fig. 4(a), and shows the MAE versus the percentage data missing. The relatively low MAE values demonstrate that an MPS can effectively learn the distribution underlying a finite number of time series, and using the learned distribution, can accurately impute missing values. As expected, imputation performance generally increased with the number of observed data points, as the MPS can leverage more conditioning information when the proportion of time-series samples that are observed increases. This effect does tend to saturate, however, as once the MPS gains enough information to determine the phase of the signal, apart from noise, it can then accurately approximate all other values. Six selected examples of the MPS-imputed time series for varying percentages of missing data are shown in subplots Figs 4(a)(ii)-(vii). Here, we also consider examples with multiple windows of missing data, which are not included in the evaluation presented in Fig. 4(a). In all cases, MPSTime could accurately infer the missing values. We observed that increasing the bond dimension $\chi_{\rm max}$ led to better imputation performance, with the most expressive MPS ($\chi_{\rm max}=40$) outperforming the 1-NNI baseline across all percentages of missing data, and for both noise levels. On the other hand, $\chi_{\rm max}=30$ performed similarly to $\chi_{\rm max}=40$ indicating that $\chi_{\rm max}=30$ is sufficiently expressive to learn the features of this dataset. We note that although there was an expected increase in MAE when increasing the noise level from $\sigma=0.1$ and $\sigma=0.2$, the $\chi_{\rm max}=30$ and $\chi_{\rm max}=40$ MPS could still accurately impute missing data when $<85\%$ of the data was missing. The imputation results here indicate that, for these bond dimensions, the MPS is sufficiently expressive to learn the joint probability distribution of the noisy sinusoid time series and can accurately impute missing values even when only a fraction of time-series values are observed.
Having demonstrated that our MPS framework can successfully encode the joint distribution underlying the simple trending sinusoidal dataset, we next aimed to test the ability of the MPS to model more complex synthetic time-series processes. To achieve this, we now use the more challenging setting and investigate several iterations of complexity by allowing the trend $m$ and/or the period $\tau$ to vary while holding the noise fixed at $\sigma=0.1$. We used the same train-test ratio as the simple dataset, however, this time training a $d=20,\chi_{\rm max}=150$ MPS for 10 sweeps. The larger value of $d$ was chosen to improve the resolution of the encoding for imputation, while the larger $\chi_{\rm max}$ value was chosen to allow the model to express the more complex correlations of the more challenging dataset. Figure 4(b) shows the mean imputation error versus the percentage of data missing across the four datasets, each with a different number of trends and periods. For almost all values, the MPS is able to outperform the 1-NNI baseline. We observe that for $<85\%$ of the data missing, the value of the MAE becomes largely fixed. This is similar to the simple dataset where, once a sufficient amount of the data is present, the projected MPS can determine the phase, period, and trend of the sinusoid and successfully fill in missing values. We note that the MPS performed best on the most complicated dataset, which contained three trends and periods. This is likely because of the large bond dimension $\chi_{\rm max}=150$ which is sufficiently expressive to learn the features of complex datasets, but can lead to overfitting on simpler ones. In addition to this, six selected examples for varying percentages of missing data are shown in subplots Figs 4(b)(ii)-(vii). All examples indicate successful imputation of missing data in the more challenging dataset.
@@
In general, the results presented in Figs 4(a)-(b) demonstrate that the MPS framework can learn the joint probability distribution of a time-series process to a degree such that it can accurately infer the probabilistic model of trending sinusoidal functions. The successful imputation of the more complicated dataset indicates that the MPS is sufficiently expressive to encode the generative model of a process with varying period, trend and phase. This first successful result shows that MPS has significant potential as a time-series analysis tool for generative modeling of a broad class of datasets with complex correlation structures.
@@ III.2.2 Real-world time series
Building on our demonstration of the imputation capabilities of the MPS, given synthetic time series from a known generative process, we now investigate its applicability to real-world time-series datasets. For our imputation tasks, we focused on time-series datasets from three domains of application to highlight the broad utility of our approach to: (i) medical data (the ‘ECG’ dataset); (ii) industrial data (the ‘Power Demand’ dataset), and (iii) astronomy (the ‘Astronomy’ dataset). To ensure that our imputation results were not biased by any particular train–test split, we adopted the dataset resampling strategy for time-series classification tasks in Bagnall et al. [60]. Specifically, we generated 30 resampled train-test splits for each dataset, then for each split we trained an MPS on the training set and evaluated its imputation error on instances from the test set, across various percentages of missing data. For any given percentage data missing, we imputed up to 15 randomly selected missing data block locations to account for the possibility that some regions may be easier to impute than others. Using the same folds and window locations for each dataset, we repeated the imputation experiment with the four baselines: 1-NNI, CSDI, BRITS, and CDRec. For demonstration purposes, we chose MPS hyperparameters that performed well on the first train-test split of each dataset based on a very simple grid search, and then held these values constant on the ‘unseen’ splits. For ECG and Power Demand, we trained a $d=10$, $\chi_{\rm max}=20$ MPS for 5 sweeps, and on the Astronomy dataset, we trained a $d=12$, $\chi_{\rm max}=35$ MPS for 3 sweeps. For the ECG and Power Demand datasets, we measured imputation performance with MAE, while the normalized MAE (NMAE) was used for the Astronomy dataset (due to differences in scale imposed by the individual windowing).
@@
Figure 5: MPSTime demonstrates competitive and often superior performance for time-series imputation on real-world datasets. The window location and test set-averaged MAE and 95% CI across 30 folds is shown in subplot (i) as a function of percentage data missing for each dataset: (a) ECG, (b) Power Demand, (c) Astronomy. For ECG and Power Demand, we trained a $d=10$, $\chi_{\rm max}=20$ MPS, and for the Astronomy dataset, we trained a $d=12,\chi_{\rm max}=35$ MPS. For each dataset, in panels (ii-v), we show representative examples of the MPS-imputed time-series (solid red line) and uncertainty due to encoding error (shaded ribbons) alongside the 1-NNI baseline (solid gold line). The shaded windows correspond to segments of observed time-series values, while the transparent windows correspond to missing data blocks. The ground-truth (unobserved) time-series values (gray line) are plotted for comparison with the imputed values.
The imputation performance of the MPS relative to the various baselines is shown in Fig. 5(a)-(c) for the ECG, Power Demand, and Astronomy datasets respectively. Starting with the ECG dataset, Fig. 5(a)(i) shows the mean MAE versus the percentage of data missing in the imputation. We observe that the MPS has superior performance to other baselines for all values of data missing. As with the synthetic data, once a sufficient number of data points are present, in this case $<85\%$ missing, the performance of the MPS maximizes and saturates. Again, this indicates that with MPS is able to lock on to the signal once a sufficient amount of information is present. Figures 5(a)(ii)–(v) show selected examples of imputation performed for different window sizes and positions in the ECG dataset. Here, we emphasize that the functional form of the ECG dataset is highly nonlinear and more complex than the synthetic datasets presented in the previous subsection. Nevertheless, the MPS framework can still learn a joint probability distribution of the dataset to a sufficient degree that it can accurately impute missing data points with only a fraction of data being present.
@@
Figure 5(b)(i) shows the imputation results for the Power Demand dataset. This dataset is shorter than previous examples and with 24 samples in each series. Again, the MPS showed superior performance to all other methods tested, and could accurately impute missing points. As with previous results, the performance of the MPS tended to saturate when $<85\%$ of the data was missing, again indicating that the MPS framework can learn the trend of the time series with very few conditioning points. This is in contrast to other methods such as CSDI and BRITS whose performance was competitive with the MPS when $5\%$ of the data was missing, but degraded significantly when larger fractions of the data were missing. Figure 5(b)(ii)-(v) show four selected examples of imputation in the Power Demand dataset with different configurations of missing data.
Finally, Fig. 5(c)(i) shows the imputation performance for the Astronomy dataset. Here the performance is split between the RR-Lyrae variable instances and the non variable instances. This is because, as seen in the examples in Fig. 5(c)(ii)-(v) these two sets of time series have very different properties with the RR-Lyrae variable being almost perfectly periodic, while the non variable time series are not. Here, the 1-NNI baseline showed superior performance for the RR-Lyrae variable data. This is not unexpected due to the periodic nature of the time series in this dataset. Since the 1-NNI method finds a sample from the training data that is closest to the data that is being imputed, for periodic datasets, there is a high probability that the training data contains a time series that is very close to the one being imputed. On the non-variable dataset CDRec outperformed MPSTime, although here, MPS performed better than CSDI and 1-NNI. Unlike the previous examples, the performance of the MPS did not show a significant performance improvement with less data missing, or a saturation effect after a fixed number of data points were known. This was particularly visible for the non variable dataset, which showed approximately constant MAE for different fractions of data missing.
@@
The imputation results presented for real-world datasets in Fig. 5 show that MPSTime can learn joint probability distributions of challenging, highly nonlinear functions that arise in real-world time series. A notable aspect of the MPS performance was that, for the ECG and Power Demand datasets, the joint probability distribution was learned with a sufficient degree of accuracy that when $<85\%$ of the data was missing, MPSTime could perform imputation with a MAE value of $<0.3$ for ECG and $<0.2$ for Power Demand. For these datasets, MPSTime outperformed all other methods. Although the MPS method did not outperform in the Kepler dataset, overall, its performance was among the best in all the imputation tests we performed. The results in Fig. 5 highlight that we can learn complex joint probability distributions that underlie real-world time-series processes well enough to perform accurate imputation. MPSTime therefore offers a very promising approach for performing imputation on a diverse array of real-world time-series datasets for a range of applications.
@@ III.3 Time-series classification
Having shown that MPSTime can accurately impute missing values in a range of challenging scenarios, including those involving real-world datasets, we now demonstrate its ability to infer data classes through a series of time-series classification (TSC) tasks. To underscore the inherent versatility of MPSTime in addressing diverse ML objectives within a unified framework, we constructed classification tasks from the same open datasets that seeded the real-world imputation experiments in Sec. III.2 and detailed in Sec. III.1.
@@
For each TSC task, we compare the performance of MPSTime to three baselines, each selected to represent a distinct algorithmic paradigm within the TSC literature. Our selection of baselines comprises: (i) Nearest Neighbor with Dynamic Time Warping (1-NN-DTW) [60]; (ii) InceptionTime [61]; and (iii) HIVE-COTE V2.0 (HC2) [62]. The classical elastic distance-based classifier, 1-NN-DTW, is a standard benchmark against which many new algorithms are compared, given its strong performance across various TSC problems on the UCR repository [60]. At the time of constructing our classification tasks, InceptionTime was considered best in category for deep learning [63]. While a recent variant, H-InceptionTime [64], has shown small but significant improvements in classification accuracy over the original InceptionTime algorithm [63], to the best of our knowledge, there are no readily available implementations for this variant. Finally, HC2 is a meta-ensemble of classifiers, each built on different time-series data representations (e.g., shapelets, bag-of-words based dictionaries, among others). Recent benchmarking studies have placed HC2 as the current top-performing algorithm for TSC, evaluated across all datasets in the UCR repository [63]. Further discussion about the TSC baselines is provided in Appendix G. We emphasize that our selection of HC2 and InceptionTime as TSC baselines – both highly expressive and computationally intensive – establishes an especially challenging setting in which to compare the performance of MPSTime.
To benchmark MPSTime against the various baselines, we evaluated classification accuracy using cross-validation across 30 stratified train-test resamples, as per Middlehurst et al. [63]. For the ECG and Power Demand datasets, we used the original train-test split published on the UCR repository as the first fold, and generated 29 additional folds using reproducible random seeds. For these datasets, we report the published 30-fold classification accuracy results for the baselines (available at the UCR repository [52]).
As the Astronomy dataset was derived as a two-class subset of the standard KeplerLightCurves open dataset (see Sec. III.1.2), published results are not publicly available for the baselines. Therefore, we evaluated reference implementations of 1NN-DTW, InceptionTime and HC2 on the Astronomy dataset, following the same 30-fold train-test resample approach outlined above. Training details, including hyperparameter settings and specific implementations for the baselines, are in Appendix G. For MPSTime, we chose parameters that performed well in our preliminary experiments on the Astronomy dataset. We note that this particular choice does not confer an unfair advantage: the 30-fold resampling methodology safeguards against configuration bias, as superior performance must be demonstrated consistently across multiple evaluation contexts (i.e., different train-test splits), rather than relying on fold-specific optimizations that may not generalize.
The classification results, shown in Fig. 6, highlight the strong performance of MPSTime, which achieves TSC accuracies competitive with state-of-the-art ML methods on our benchmarks. Notably, for both ECG and Power Demand datasets, MPSTime achieves accuracies that closely match (within statistical uncertainty) the top-performing deep learning and ensemble methods, InceptionTime and HC2, respectively, while outperforming the classic 1-NN-DTW benchmark on these benchmarks. On the Astronomy dataset, we observed the greatest variability in performance differences across different methods, with MPSTime achieving a mean accuracy of $0.91\pm 0.01$, which was lower than both InceptionTime ($0.93\pm 0.01$) and HC2 ($0.96\pm 0.01$). It is worth noting that with the Astronomy TSC task, we did not perform extensive hyperparameter tuning for MPSTime as for ECG and Power Demand. The baseline methods, while evaluated with default parameters, benefit from extensive prior optimization studies that inform these default configurations. A systematic exploration of MPSTime’s parameter space, comparable to those conducted for the baseline methods, could yield more competitive results for the Astronomy dataset.
[…]
@@
Figure 6: MPSTime outperforms a classic time-series classification benchmark (1-NN-DTW) and is competitive with modern state-of-the-art algorithms. The mean classification accuracy and 95% CI (across cross-validation folds) are shown for each of the three empirical datasets: Power Demand, ECG, and the Astronomy dataset. We compare our MPS classifier with three baselines: one nearest neighbor with Dynamic Time-Warping (1-NN-DTW), InceptionTime, and HIVE-COTE 2.0 (HC2).

### v2

@@ III.2.1 Synthetic time series
In this section we investigate the ability of MPSTime to impute missing values from time-series data distributions of noisy trendy sinusoids (NTS) with incrementally increasing diversity, generated through different choices of the parameters in the common generative model in Eq. (22). For our simplest imputation setting, NTS1, we trained MPS with three different fixed bond dimensions, $\chi_{\rm max}=20,30,40$ for ten sweeps. The learning rate $\eta$ and physical dimension $d$ were both determined by hyperparameter tuning with search ranges summarized in Appendix C.1. For a given percentage of missing data in each unseen test instance, we masked a random contiguous block of samples, and the missing segment was imputed using MPSTime.
Our first key result is presented in Fig. 4(a), and shows the imputation error, reported as the mean absolute error (MAE), versus the percentage of missing data. The relatively low MAE values for all bond dimensions $\chi_{\rm max}$ demonstrate that an MPS can effectively learn the distribution underlying a time-series dataset, and using the learned distribution, accurately impute missing values. This is reflected in the improvement in imputation performance as the bond dimension was increased, as more expressive MPS are able to capture more correlations, and therefore posses the capacity to infer missing values more accurately. For all $\chi_{max}$ values, imputation performance improved as the number of observed data points increased, this is expected, as less missing data provides the MPS with additional conditioning information. This improvement tends to saturate for $<45\%$ missing data, as once the MPS gains enough information from the observed points to infer the phase of the signal to a sufficient degree of accuracy, aside from noise, it can then approximate all other values to a similar level of accuracy.
We observed a similar saturation in performance for the 1-NNI benchmark (dashed black curve), however, occurring at a higher MAE. This is because only a small number of observations are required to fix the nearest training example, after which 1-NNI repeatedly returns the same example, and therefore cannot reduce its error below the residual train–test instance mismatch (within noise).
To demonstrate the imputation performance of MPSTime visually, four selected examples of the MPSTime-imputed time series for varying percentages of missing data are shown in Figs 4(a)(ii)-(v). Fig. 4(a)(v) shows an extension of the imputation task to a multi-window setting (note that for simplicity, multi-window imputation data is not included in Fig. 4(a)(i)], however we expect similar imputation performance to the single-window setting).
All three MPS models with bond dimensions $\chi_{\rm max}=20,30,40$, shown as the blue, orange and green curves, respectively, in Fig. 4(a) consistently outperformed the 1-NNI benchmark (dashed black curve) across all missing data percentages. The imputation results here indicate that, for small to moderate bond dimensions of $\chi_{\rm max}$ = 20-40, the MPS is sufficiently expressive to learn the joint probability distribution of the noisy sinusoid time series with fixed trend and unknown phase, and can accurately impute missing values, achieving MAE below 0.1 even when 85% of time-series data points are missing.
Having demonstrated that MPSTime can successfully encode the joint distribution underlying a relatively simple noisy periodic dataset, we next aimed to test its ability to model a more complex range of oscillatory dynamics. To achieve this, we investigated four more complex versions of the underlying process in Eq. (22) (NTS2–5) by allowing the trend $m$ and/or the period $\tau$ to vary while holding the noise scale fixed at $\sigma=0.1$. The optimal values for $\chi_{\rm max}$, $\eta$ and $d$ were determined via a random search, as described in Appendix C.1.
Figure 4(b) shows the mean imputation error versus the percentage of data missing across the four datasets (NTS2–5). We observe that the MAE for 1-NNI and MPSTime has already approximately saturated when the proportion of missing data has dropped to around $85\%$. This is similar to the simple dataset (NTS1) in Fig. 4(a) where, once a sufficient amount of the data is present, the projected MPS can accurately infer the relevant unknown properties of the process (e.g., the phase $\psi$, period $T$, and/or trend $m$), and accurately impute missing values.
Across the challenging datasets in Fig. 4(b)(i), MPSTime imputation performance (solid curves) decreased with dataset diversity, reflecting the increasingly difficult challenge of modeling a broader range of temporal correlation structures. As shown in Fig. 4(b)(i) MPSTime performed best on the two datasets with least diversity, NTS2 with three possible periods and fixed trend (solid blue curve) and NTS3 with three possible trends and fixed period (solid orange curve). With the exception of NTS5 (solid red curve), MPSTime outperformed the 1-NNI baseline (dashed curve for each dataset) for almost all percentages of missing data. Our hyperparameter search for NTS5 consistently selected the largest possible bond dimension $\chi_{\rm max}=160$, suggesting that further increasing the expressive capacity of the MPS through larger $\chi_{\rm max}$ could yield further improvements, particularly for $>45\%$ data missing. As we will see in the next section, the nearest-neighbor benchmark is also unexpectedly effective at imputing complex univariate data, so achieving parity with our benchmark is not an indicator of poor performance.
Four examples depicting the imputation performance for varying percentages of missing data in the most challenging setting (NTS5), with three distinct periods, trends, and random phase, are shown in Figs 4(b)(ii)–(v). The first three subplots, Figs 4(b)(ii)–(iv), correspond directly to the analysis presented in Fig. 4(b)(i), where performance of MPSTime is evaluated for a single contiguous block of missing data positioned at various locations within the time-series instances. The last subplot, Fig. 4(b)(v), showcases MPSTime’s imputation behavior when multiple missing periods are present. While this scenario is not explicitly considered in the analysis presented in Fig. 4(b)(i), we include it here as visually compelling example to highlight the robustness and flexibility of MPSTime across diverse missing-data settings.
@@
In general, the results presented in Figs 4(a),(b) demonstrate that the MPS framework, when given sufficient training data, can learn the joint probability distribution of a dynamical process to a degree such that it can accurately infer the probabilistic model underlying trending sinusoidal functions. The successful imputation of the more complicated dataset, NTS5, indicates that the MPS is sufficiently expressive to encode the generative model of a process with varying period, trend, and phase. Together, these results support the usefulness of MPS as the basis for developing novel algorithms for analyzing time series.
@@ III.2.2 Real-world time series
Having demonstrated that MPSTime accurately imputes missing values from unseen time series, we now investigate its performance on real-world time-series datasets from three domains of application: (i) medicine (the ‘ECG’ dataset); (ii) industry (the ‘Power Demand’ dataset); and (iii) astronomy (the ‘Astronomy’ dataset). Further details about these datasets appear in Sec. III.1.2. To avoid bias from any specific train–test split, we used a cross-validation strategy with 30 resampled train–test splits for each dataset [66]. We tuned a separate MPS on each train–test split using 5-fold cross validation on the training set, then evaluated imputation error on unseen test instances. Hyperparameter optimization was performed using a random search variant, with tuning ranges provided in Appendix C.1. For each test instance, we evaluated imputation performance by removing a contiguous segment of values, equivalent to a percentage of the total time-series length. As in Sec. III.2.1, we examined missing data scenarios from 5% to 95%, in 10 percentage point increments. At each percentage level, we randomly varied the starting position of the missing segment to prevent bias from temporal segments with potentially more predictable patterns than others. Using the same train–test splits and missing segment locations, we evaluated the imputation performance of the four benchmark algorithms described in Sec. III.2: 1-nearest-neighbor imputation (1-NNI), a generative model-based imputation approach (CSDI), a recurrent neural network (BRITS-I), and a matrix completion method (CDRec).
@@
Figure 5: MPSTime exhibits competitive (and often superior) performance to specialist benchmark algorithms for time-series imputation on real-world datasets. Imputation error, reported as the Mean Absolute Error (MAE) and 95% CI across 30 test folds is shown as a function of percentage data missing for each dataset in panels (i): (a)(i) ECG, (b)(i) Power Demand, (c)(i, iv) Astronomy (see Sec. III.1.2). MPSTime (red) is compared against four benchmarks: CDRec (blue), CSDI (green), 1-NNI (gold), and BRITS-I (purple). For Astronomy, the imputation performance on (c)(i) $\gamma$ Doradus and (c)(iv) RR Lyrae is plotted separately due to differing MAE scales. Benchmarks with substantially higher errors for $\gamma$ Doradus stars (BRITS-I and CSDI) and RR Lyrae stars (BRITS-I) were excluded from the Astronomy panels to maintain visual clarity among the top-performing algorithms. Panels (ii)–(v) for ECG and Power Demand show representative MPSTime imputations (solid red line), uncertainty due to encoding error (shaded ribbons), and the corresponding behavior of the 1-nearest neighbor (1-NNI) benchmark (solid gold line). For Astronomy, the panels (c)(ii, iii) and (c)(v, vi) show selected examples of the best performing MPSTime imputations for $\gamma$ Doradus stars and RR Lyrae stars, respectively. In each imputation example, shaded regions denote observed segments, and transparent regions indicate missing (unobserved) data blocks. Ground-truth (unobserved) time-series values (gray line) are overlaid for reference.
[…]
@@
MPSTime also exhibited superior imputation performance on the Power Demand dataset for almost all percentages of data missing, as shown in Fig. 5(b)(i). This dataset contains time series of fewer samples ($24$ samples each), compared to ECG ($96$ samples each), and MPSTime (red) exhibited a similar saturation in performance, this time at around $<55\%$ data missing, again indicating that the MPS framework can learn the complex patterns in time series with relatively few conditioning points. Among the benchmark algorithms, 1-NNI (gold) demonstrated the strongest performance across most fractions of missing data. Other benchmark imputation methods, such as CDRec (blue) and BRITS-I (purple), exhibited competitive performance with MPSTime at low % data missing (e.g., at $5\%$), but degraded significantly when larger fractions of the data were missing. Figures 5(b)(ii)–(v) show selected examples of imputation in the Power Demand dataset with different configurations of missing data, visually demonstrating the ability of MPSTime (red) to accurately impute complex nonlinear patterns.
Finally, imputation performance on the Astronomy dataset is plotted in Fig. 5(c), where the mean absolute error (MAE) is shown separately for the two classes, (i) $\gamma$ Doradus stars and (iv) RR Lyrae stars, due to several orders of magnitude difference in their respective data and error scales. For $\gamma$ Doradus stars in Fig. 5(c)(i), the matrix decomposition benchmark CDRec (blue) outperformed both MPSTime (red) and 1-NNI (gold) for $<85\%$ data missing. Unlike the previous examples, the imputation performance of MPSTime continued to improve as additional data points were made available. The two neural network-based benchmarks, BRITS-I and CSDI, were excluded from the analysis presented in Fig. 5(c)(i) due to substantially larger errors (MAE $>0.007$ for BRITS-I and MAE $>0.004$ for CSDI, for all % missing data). Selected examples of the best performing MPSTime imputations (red) are shown in Figs 5(c)(ii),(iii).
For RR Lyrae in Fig. 5(c)(iv), MPSTime (red) and 1-NNI (gold) performed similarly for $<75\%$ missing data, with both outperforming the specialist algorithms for imputation across the full % missing range. Notably, for $<95\%$ data missing, the performance of both MPSTime and 1-NNI did not show a continual improvement or saturation effect after a fixed number of data points were known. This is not unexpected due to the highly periodic nature of RR Lyrae star time series, as shown in Figs 5(c)(v),(vi), where, similar to the NTS datasets in Sec. III.2.1, MPSTime is able to infer the phase of the signal with relatively few points, beyond which there are no substantial performance gains. This performance saturation closely matches the expected behavior of 1-NNI (gold) on a highly regular and periodic dataset where matching to a near-identical example with similar phase from the training set is likely.
@@
The imputation results presented for real-world datasets in Fig. 5 show that MPSTime can learn the joint probability distributions that underlie real-world time-series datasets with sufficient precision to perform accurate imputation. A notable aspect of the MPSTime performance was that, for the ECG and Power Demand datasets, the joint probability distribution was learned with a sufficient degree of accuracy that when $<85\%$ of the data was missing, MPSTime could perform imputation with a MAE value of $<0.3$ for ECG and $<0.2$ for Power Demand. For these datasets, MPSTime almost always outperformed all other benchmarks, including specialist algorithms for time-series imputation. Although MPSTime did not outperform in the Astronomy dataset, overall, its performance was among the best in all the imputation tests we performed. MPSTime therefore offers a promising approach for performing imputation on a diverse range of real-world time-series datasets for which the underlying generative process may be unknown.
@@ III.3 Time-series classification
Having shown that MPSTime can accurately impute missing values across multiple challenging scenarios involving synthetic and real-world data, we now demonstrate its ability to infer data classes through a series of time-series classification (TSC) tasks. To underscore the inherent versatility of MPSTime in addressing diverse ML objectives within a unified framework, we constructed classification tasks from the same open real-world datasets (Sec. III.1) analyzed in the imputation experiments above.
@@
For each TSC task, we compare the performance of MPSTime to three benchmarks, each selected to represent a distinct algorithmic paradigm within the TSC literature. Our selection of benchmarks comprise: (i) Nearest Neighbor with Dynamic Time Warping (1-NN-DTW) [66]; (ii) InceptionTime [67]; and (iii) HIVE-COTE V2.0 (HC2) [68]. The classical elastic distance-based classifier, 1-NN-DTW, is a standard benchmark against which many new algorithms are compared, given its strong performance across various TSC problems on the UCR repository [66]. At the time of the analysis, the specialized deep neural-network benchmark, InceptionTime, was considered a state-of-the-art deep learning algorithm based on recent benchmarking [69]. Finally, HC2 is a meta-ensemble of classifiers, each built on different time-series data representations (e.g., shapelets, bag-of-words based dictionaries, among others). Recent results placed HC2 as the top-performing algorithm on the UCR TSC repository [69]. Further discussion about these benchmark algorithms for TSC is in Appendix G. We emphasize that our selection of HC2 and InceptionTime as benchmarks – both highly expressive and computationally intensive – establishes an especially challenging setting in which to compare the performance of MPSTime.
We evaluated classification accuracy using cross-validation across 30 stratified train-test resamples, as per Middlehurst et al. [69]. For the ECG and Power Demand datasets, we used the same train–test split permutations from the UCR repository [55], matching those used for the published 30-fold mean accuracy results. The Astronomy dataset was newly derived for this study from the KeplerLightCurves open dataset (see Sec. III.1.2). As published results are not available for the benchmarks, we evaluated reference implementations of 1-NN-DTW, InceptionTime and HC2 on the Astronomy dataset, following the same 30-fold train-test resample approach outlined above. Training details, including hyperparameter settings and specific implementations for the benchmarks are in Appendix G.
@@
Figure 6: MPSTime outperforms a classic time-series classification benchmark (1-NN-DTW) and is competitive with modern state-of-the-art classification algorithms: the neural-network-based InceptionTime and the meta-ensemble HIVE-COTE V2.0 (HC2). The 30-fold mean classification accuracy and 95% confidence interval (CI) are shown for each of the three empirical datasets: Power Demand, ECG, and Astronomy (see Sec. III.1). We compare our MPS classifier, MPSTime (red), with three benchmarks: one-nearest neighbor with Dynamic Time-Warping (1-NN-DTW) (gold), InceptionTime (green), and HIVE-COTE V2.0 (HC2) (purple).
The classification results, shown in Fig. 6, highlight the strong performance of MPSTime, which achieves TSC accuracies competitive with state-of-the-art TSC algorithms on our benchmarks. Notably, for both ECG and Power Demand datasets, MPSTime achieves accuracies that are competitive with top-performing deep learning and ensemble methods, InceptionTime and HC2, respectively, while outperforming the classic 1-NN-DTW benchmark on these datasets. On ECG, our preliminary findings (see Appendix. H) also suggest that MPSTime achieves a favorable accuracy/train-time trade-off relative to the benchmarks. Finally, on the Astronomy dataset, we observed the greatest variability in performance differences across methods, with MPSTime achieving a mean accuracy of $88.7\%$ ($95\%$ CI: $87.6\%-89.8\%$), which was lower than both InceptionTime, $91.3\%$ ($95\%$ CI: $90.2\%-92.4\%$), and HC2, $95.9\%$ ($95\%$ CI: $95.2\%-96.6\%$).
In summary, our findings indicate that an initial adaption of MPSTime to classification, without sophisticated hyperparameter tuning procedures, nevertheless achieves competitive performance (and relatively efficient training cost) on real-world time-series datasets when compared to state-of-the-art methods. These results demonstrate both the versatility of our framework in tackling multiple time-series ML objectives (i.e., both imputation and classification) and reveal significant opportunities for improving the performance of MPSTime through further specialization.